OpenStudy (anonymous):

Consider the solid consisting of the first octant portion of a sphere of radius 4, centered at the origin, and lying below the plane z=2. find the volume of this solid as a triple integral.

ganeshie8 (ganeshie8):

try \[\dfrac{1}{8} \frac{4}{3}\pi 4^3 - \iiint 1. dV\]

OpenStudy (anonymous):

can we do this without volume formula, only with triple integral?

OpenStudy (anonymous):

@ikram002p

OpenStudy (anonymous):

could you also please express the whole integral with boundaries?

OpenStudy (anonymous):

latex is not showing up on my machien :O http://prntscr.com/52lsei

OpenStudy (ikram002p):

logout close ur browser login from other account xD

OpenStudy (anonymous):

you can only type boundaries and integrate, can you? @beccaboo022a

OpenStudy (ikram002p):



OpenStudy (anonymous):